OpenStudy (anonymous):

Find the 9th term of the arithmetic sequence. a7=-42 a11=-28

OpenStudy (imstuck):

The first term, a1 is 7. The formula for an arithmetic sequence is \[a _{n}=a _{1}+(n-1)d\]. You need to write to equations and solve them simultaneously for a1 and d, then rewrite the equations. Lets see how this works out. The first equation is -42 = a1 + (7-1) d; the second one is -28 = a1 + (11-1) d. We could subtract the 1 from the 7 and that's 6, and then 11-1 = 10. Now we have -42 = a1 + 6d and -28 = a1 + 10d. Multiply the first equation by -1 to rid yourself of the a1. Now we have 42 = -a1 - 6d and -28 = a1 + 10d. The a1's cancel out, leaving 14 = 4d. Divide by 4 and get d = 3.5. Now fill in the 3.5 in one of the equations to find a1. -42 = a1 + 6(3.5). That gives us -63 is a1, which makes sense, since the 7th term is -42 and the 11th term is -28. Got that?